# On $\operatorname{CAT}(\kappa)$ surfaces

### v2

@@ 4. Vertex-Edge Triangulations
In this section, we define vertex-edge triangulations and establish their main properties in preparation for the first proof of 1.1. Throughout this section, $X$ is a $\operatorname{CAT}(\kappa)$ surface.
@@ 4.1. Existence of vertex-edge refinements
First, we have a lemma about recognizing vertex-edge triangulations.
@@
Let $P$ be a polygon in a $\operatorname{CAT}(\kappa)$ surface of perimeter less than $2D_{\kappa}$ such that $P$ is the union of non-overlapping convex polygons $P_{1},\ldots,P_{n}$, where each $P_{i}$ is a triangle or quadrilateral. Assume that each polygon $P_{i}$ intersects $P_{i+1}$ along a single common edge, and that the $P_{i}$ do not intersect otherwise except at vertices. Moreover, if $P_{i}$ is a quadrilateral, then we assume that it does not intersect $P_{i-1}$ and $P_{i+1}$ along adjacent edges of $P_{i}$. Finally, we require each original vertex of $P$ to be a vertex of one of the $P_{i}$. We form a triangulation $\mathcal{T}$ of $P$ by dividing each $P_{i}$ that is a quadrilateral into triangles by connecting two non-adjacent vertices with a geodesic. Then $\mathcal{T}$ is a vertex-edge triangulation of $P$.
@@
The proof is somewhat technical, but the geometric idea at its core can be easily read off from Figure 4.1
@@
    For each vertex $v_{j}$ not contained in $[pq]$ or $[pr]$, extend the unique geodesic $[pv_{j}]$ to a geodesic that intersects the opposite edge $[qr]$ using 3.3. See Figure 1(b). This yields a collection of distinct geodesics $\{I_{k}\}_{k=1}^{K}$, enumerated in increasing order based upon where the endpoint lies on $[qr]$, oriented from $q$ to $r$. Note that two geodesics $I_{j_{1}},I_{j_{2}}$ cannot share the same endpoint, since otherwise they must coincide by uniqueness of geodesics. Also, set $I_{0}=[pq]$ and $I_{K+1}=[pr]$. For all $k\in\{0,\ldots,K\}$, the geodesics $I_{k},I_{k+1}$ together with a subarc of $[qr]$ bound a triangle $T_{k}$.
@@
    For each triangle $T_{k}$ apply the following. Let $\mathcal{A}_{k}$ denote the collection of polygons of the form $A_{i}\cap T_{k}$ intersecting the interior of $T_{k}$; denote this polygon by $T_{k}^{i}$. By 2.3, $T_{k}^{i}$ is a convex polygon. Note that each vertex of each $T_{k}^{i}$ is contained in $I_{k}\cup I_{k+1}$. This implies that every edge of some $T_{k}^{i}$ not contained in $I_{k}$ or $I_{k+1}$ must connect $I_{k}$ to $I_{k+1}$. Enumerate such distinct edges as $\{J_{k}^{j}\}_{j=1}^{J_{k}}$ in increasing order from $[qr]$ to the vertex $p$. Also, denote the subarc of $[qr]$ contained in $T_{k}$ by $J_{k}^{0}$. Note that $J_{k}^{j}$ and $J_{k}^{j+1}$ along with the relevant subarcs of $I_{k}$ and $I_{k+1}$ form either a triangle or quadrilateral. We subdivide each triangle $T_{k}$ as follows. For all $j\in\{0,\ldots,J_{k}-1\}$, if $J_{k}^{j}$ and $J_{k}^{j+1}$ do not share an endpoint, then connect an endpoint of $J_{k}^{j}$ to an endpoint of $J_{k}^{j+1}$ by a geodesic, forming two triangles. In this manner we obtain a triangulation $\mathcal{T}$ of $P$. The final product is shown in Figure 1(c).
@@
It follows from construction that $\mathcal{T}$ is a vertex-edge triangulation of $T$. Indeed, $\mathcal{T}$ is obtained first by drawing geodesics from $p$ to the opposite edge to form the subtriangles $T_{k}$. Then each $T_{k}$ is the union of triangles satisfying the requirements of 4.3. We conclude that $\mathcal{T}$ is vertex-edge.
The final property is that, for each $A\in\mathcal{A}$, the collection $\mathcal{T}_{A}$ can be obtained by vertex-edge subdivision of $A$. Observe that the geodesics $\{I_{k}\}_{k=1}^{K}$ divide $A$ into the non-overlapping union of triangles and quadrilaterals satisfying the requirements of 4.3. We now apply this lemma to conclude that $\mathcal{T}_{A}$ is a vertex-edge triangulation of $A$. ∎
@@ 4.2. Subdividing triangles
Next, to prove (2), the main task is to show that the comparison triangles $\bar{T}_{1}$ and $\bar{T}_{2}$ can be fitted into $\bar{T}$ such that $\bar{T}_{1}$ and $\bar{T}_{2}$ are non-overlapping. Our argument is based on Alexandrov’s Lemma; see [9, I.2.16]. Consider the triangle $\bar{T}_{1}$ arranged to share the side $[\bar{p}\bar{q}]\subset\bar{T}$ and have overlapping interior. Similarly, the triangle $\bar{T}_{2}$ is arranged to share the side $[\bar{p}\bar{r}]\subset\bar{T}$ and have overlapping interior. Let $\bar{y}$ denote the third vertex of $\bar{T}_{1}$ and let $\bar{z}$ denote the third vertex of $\bar{T}_{2}$. The situation is summarized in Figure 4.2.
@@
Since $\bar{\beta}_{1}\leq\bar{\beta}$ and $\alpha_{1}\leq\alpha$, we see that $\bar{y}$ is contained in $\bar{T}$. Likewise, $\bar{z}$ is also contained in $\bar{T}$. This implies that the triangles $\bar{T}_{1}$ and $\bar{T}_{2}$ are entirely contained in $\bar{T}$. Moreover, since $\bar{\alpha}\geq\bar{\alpha}_{1}+\bar{\alpha}_{2}$, the segments $[\bar{p}\bar{r}]$, $[\bar{p}\bar{z}]$, $[\bar{p}\bar{y}]$, $[\bar{p}\bar{q}]$ emanate from $\bar{p}$ counterclockwise in this order, with possibly $[\bar{p}\bar{z}]$, $[\bar{p}\bar{y}]$ coinciding. From this it follows that $\bar{T}_{1}$ and $\bar{T}_{2}$ do not overlap. We conclude that $|\bar{T}_{1}|+|\bar{T}_{2}|\leq|\bar{T}|$. ∎
@@ 5. $\operatorname{CAT}(\kappa)$ surfaces have bounded curvature
5.1. The first proof
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface and let $x\in X$. Let $r>0$ be sufficiently small so that $B(x,r)$ is $\operatorname{CAT}(\kappa)$ and convex in $X$. We assume without loss of generality that $\kappa=1$. Use 2.5 to find a triangulation $\mathcal{S}$ of $B(x,r)$ with mesh less than $2D_{\kappa}$. Let $S_{1},\ldots,S_{n}$ be the collection of triangles in $\mathcal{S}$ that contain $x$. Then $S=\bigcup_{i=1}^{n}S_{i}$ is a closed topological disk such that $x\in S^{\circ}$. This follows from the convexity of each triangle and uniqueness of geodesics. We claim that $S^{\circ}$ is the neighborhood of $x$ as required by the definition of surface of bounded curvature.
Let $\mathcal{T}=\{T_{j}\}_{j=1}^{m}$ be an arbitrary collection of simple non-overlapping triangles in $S^{\circ}$. Our goal is to find a uniform upper bound on $\delta(\mathcal{T})$. We apply 4.4 to find for each $T_{j}\in\mathcal{T}$ a triangulation $\mathcal{T}_{j}$ of $T_{j}$ that is a vertex-edge refinement of the collection of polygons $S_{i}\cap T_{j}$, $i=1,\ldots,n$, intersecting the interior of $T_{j}$ with respect to $T_{j}$. Note that 2.3 implies that each set $S_{i}\cap T_{j}$ is in fact a polygon provided that its interior is non-empty. Observe from 4.5 that $\delta(T_{j})\leq\delta(\mathcal{T}_{j})$.
Next, for each triangle $S_{i}$, let $\mathcal{S}_{i}$ be the set of triangles in $\bigcup_{j}\mathcal{T}_{j}$ contained in $S_{i}$. We apply 4.4 a second time to $\mathcal{S}_{i}$ to find a vertex-edge refinement of $\mathcal{S}_{i}$ with respect to $S_{i}$. Denote this by $\mathcal{S}_{i}^{\prime}$. Suppose that $T\in\mathcal{T}_{j}$ and contained in $S_{i}$. Let $\mathcal{S}_{i}^{\prime}(T)$ denote the subset of $\mathcal{S}_{i}^{\prime}$ contained in $T$. Then, according to 4.4, $\mathcal{S}_{i}^{\prime}(T)$ is a vertex-edge triangulation of $T$ and hence $\delta(T)\leq\delta(\mathcal{S}_{i}^{\prime}(T))$ by 4.5. This gives
@@ 5.1. The first proof
Here, $\bar{T}_{1}^{\prime}$ is the model triangle in $M_{1}^{2}$ corresponding to $T^{\prime}$. Similarly, for each triangle $S_{i}$ we let $\bar{S}_{i}$ denote the model triangle of $S_{i}$ in $M_{1}^{2}$. Since $\mathcal{S}_{i}^{\prime}$ is a vertex-edge refinement with respect to $S_{i}$, 4.5 gives
@@
5.2. The second proof
Let $(X,d)$ be a $\operatorname{CAT}(\kappa)$ surface. Without loss of generality, we assume that $\kappa=1$. Given a point $x\in X$, let $U$ be a neighborhood of $x$ with compact closure. As shown in Section 3.4, for such a neighborhood $|U|$ is finite; note that this is the only step in this proof that relies on 2.5 concerning triangulations. We may choose $U$ to be sufficiently small so that $\delta(T)\leq\pi/20$ for any triangle $T\subset U$, 3.1 holds for $\theta=\pi/20$, and the model triangles $\bar{T}_{0}$ and $\bar{T}_{1}$ are $2$-bi-Lipschitz equivalent for any triangle $T\subset U$.
To verify that $X$ has bounded curvature, it is enough to show that there is a constant $C\geq 1$ such that
(5.1) $\delta(T)\leq C|T|$
for all simple triangles $T\subset U$. Let $T$ be such a triangle. If $\delta(T)\leq 0$, then the claim holds trivially. Thus we may assume that $\delta(T)>0$. Ideally, we would be able to obtain (5.1) as a consequence of 3.1. The obstacle is that $T$ may have a small angle. The strategy of our proof is to subdivide $T$ into two parts: a collection of triangles to which 3.1 applies and at most two triangles of small diameter.
Denote the angles of $T$ by $\alpha,\beta,\gamma$, arranged so that $\alpha\geq\beta\geq\gamma$. Let $p,q,r$ denote the vertices at $\alpha,\beta,\gamma$, respectively, and $A,B,C$ the edges opposite $\alpha,\beta,\gamma$, respectively. Observe in particular that $\alpha\geq\pi/3$. In the case that $\gamma\geq\pi/20$, we apply 3.1 (3) to conclude that $|T|\geq\delta(T)/(8L^{2})$ for some fixed constant $L$. This verifies (5.1) in the special case that $\gamma\geq\pi/20$.
[…]
Figure 5.1. The triangle $T$ in the second proof, following reductions
Thus we assume in the following that $\gamma\leq\pi/20$. As a consequence, observe that $\alpha\in[9\pi/20,\pi)$ (the possibility $\alpha=\pi$ violates the uniqueness of geodesics for $\operatorname{CAT}(\kappa)$ surfaces). In the case that $\alpha\geq 14\pi/20$, we perform an initial subdivision of $T$. Let $E$ be an arc that bisects the angle $\alpha$ and intersects $A$ at a point $e$. This forms two subtriangles $T_{1}=[pqe]$ and $T_{2}=[pre]$. Let $\alpha_{1}$ denote the angle at $p$ in the triangle $T_{1}$ and $\alpha_{2}$ the same for $T_{2}$. Then $\alpha_{1},\alpha_{2}\in[7\pi/20,\pi/2)$. Note that $\beta\leq 6\pi/20$. We see that for $T_{1}$, one of the three possibilities holds: the excess $\delta(T_{1})$ is negative; each angle of $T_{1}$ is at least $\pi/20$; or, the angle $\beta$ is at most $\pi/20$. The first two cases are handled by the previous observations. In turn, for the triangle $T_{2}$, we already have that $\gamma\leq\pi/20$. Since $\delta(T)\leq\delta(T_{1})+\delta(T_{2})$, it is enough to prove (5.1) for the triangles $T_{1}$ and $T_{2}$ separately.
In this way, we reduce to the case that $T$ has largest angle $\alpha\in[9\pi/20,14\pi/20]$ and smallest angle $\gamma\leq\pi/20$. Observe as a consequence that $\beta\in[5\pi/20,12\pi/20]$. See Figure 5.1.
We apply the following procedure to split off good triangles from $T$. Assume we have a triangle $T=T_{0}^{\prime}$ such that $\alpha,\beta\in[4\pi/20,17\pi/20]$. By relabeling vertices if needed, we may assume that $\alpha\geq\beta\geq\gamma$, and in particular (given that $\delta(T)>0$) that $\alpha\geq 9\pi/20$. Let $C_{1}$ be the arc bisecting the angle $\alpha$, which intersects $A$ at a point $q_{2}$. This splits $T$ into two triangles $T_{1}=[pq_{2}q]$ and $T_{1}^{\prime}=[pq_{2}r]$. Let $\beta_{1}$ denote the angle at $q_{2}$ in $T_{1}$ and $\beta_{1}^{\prime}$ denote the angle at $q_{2}$ in $T_{1}^{\prime}$. Define $\alpha_{1},\alpha_{1}^{\prime}$ similarly. Note that $\alpha_{1}^{\prime}\geq 4\pi/20$, which implies that $\beta_{1}^{\prime}\leq 17\pi/20$. This in turn implies that $\beta_{1}\geq 3\pi/20$. Thus all the angles of $T_{1}$ are at least $203\pi/$ 20, and we can apply 3.1 to $T_{1}$. On the other hand, the triangle $T_{1}^{\prime}$ satisfies $\alpha_{1}^{\prime},\beta_{1}^{\prime}\in[4\pi/20,17\pi/20]$, the same angle conditions as $T$ itself. Let $p_{1}$ denote the vertex of $T_{1}$ lying on the edge $B$ and $q_{1}$ denote the vertex on the edge $A$. Thus we may repeat the procedure to obtain triangles $T_{2},T_{3},T_{4},T_{5},\ldots$, vertices $p_{2},q_{2},p_{3},q_{3},\ldots$, where each $p_{n}$ lies on $B$ and each $q_{n}$ lies on $A$. By 3.1, the lengths of each of the edges of $T_{n}$ are comparable.
We need to show that we can terminate this process after finitely many steps. More precisely, we claim that for all $\varepsilon>0$ one can find $n\in\mathbb{N}$ such that $\delta(T_{n})<\varepsilon$, which follows if we can show that $\operatorname{diam}(T_{n}^{\prime})\to 0$ as $n\to\infty$. First, by the comparability of the edge lengths of $T_{n}$, we can rule out the case that $p_{n}\to r$ but $q_{n}$ does not, or vice versa. Next, the same comparability rules out the case that both $p_{n}$ and $q_{n}$ converge to a limit point other than $r$. We conclude that both $p_{n}$ and $q_{n}$ converge to $r$. This implies that $\operatorname{diam}T_{n}^{\prime}\to 0$.
5.3. The third proof
Let $X$ be a locally $\operatorname{CAT}(\kappa)$ surface. Let $\mathcal{U}$ be a cover of $X$ by convex open sets of diameter less than $2D_{\kappa}$ on which the restriction of the metric is $\operatorname{CAT}(\kappa)$. Let $\varepsilon_{n}$ be a sequence of decreasing reals each less than $2D_{\kappa}$ and limiting to 00. Apply 2.5 to obtain a sequence of triangulations $\mathcal{T}_{n}$ of $X$ corresponding to the parameters $\varepsilon_{n}$ and the collection $\mathcal{U}$; in particular, each $\mathcal{T}_{n}$ has mesh less than $2D_{\kappa}$. By 4.4, we can choose the $\mathcal{T}_{n}$ so that each triangle $S\in\mathcal{T}_{n}$ for all $n\geq 2$ is contained in some triangle of $\mathcal{T}_{1}$, and so that for each $T\in\mathcal{T}_{1}$ the subset of $\mathcal{T}_{n}$ of triangles intersecting the interior of $T$ form a vertex-edge triangulation of $T$.

### v3

@@ 4. Vertex-Edge Triangulations
In this section, we define vertex-edge triangulations and establish their main properties. Throughout this section, $X$ is a $\operatorname{CAT}(\kappa)$ surface.
@@ 4.1. Existence of vertex-edge refinements
First, we have a lemma about recognizing vertex-edge triangulations. The situation in this lemma is illustrated in Figure˜4.1.
@@
Let $P$ be a convex polygon of perimeter less than $2D_{\kappa}$ that is a $\operatorname{CAT}(\kappa)$ space. Let $J_{i}$ ($i=1,\ldots,n-1$) be a set of non-crossing geodesic arcs in $P$ that intersect $\partial P$ only at the endpoints and divide $P$ into $n$ non-overlapping polygons $P_{1},\ldots,P_{n}$. Assume that each $P_{i}$ is a triangle or quadrilateral whose edges are subarcs of $J_{i-1}$ and $J_{i}$ (when these are defined) and $\partial P$. Moreover, if $P_{i}$ is a quadrilateral, then we assume that it does not intersect $P_{i-1}$ and $P_{i+1}$ along adjacent edges of $P_{i}$. Finally, we require each original vertex of $P$ to be a vertex of one of the $P_{i}$. We form a triangulation $\mathcal{T}$ of $P$ by dividing each $P_{i}$ that is a quadrilateral into triangles by connecting two non-adjacent vertices with a geodesic. Then $\mathcal{T}$ is a vertex-edge triangulation of $P$.
[…]
Figure 4.1. A polygon as in ˜4.3
@@
The proof is somewhat technical, but the geometric idea at its core can be easily read off from Figure˜4.2
@@
    For each vertex $v_{j}$ not contained in $[pq]$ or $[pr]$, extend the unique geodesic $pv_{j}]$ to a geodesic that intersects the opposite edge $[qr]$ using ˜3.1. See Figure˜2(b). This yields a collection of distinct geodesics $\{I_{k}\}_{k=1}^{K}$, enumerated in increasing order based upon where the endpoint lies on $[qr]$, oriented from $q$ to $r$. Note that two geodesics $I_{j_{1}},I_{j_{2}}$ cannot share the same endpoint, since otherwise they must coincide by uniqueness of geodesics. Also, set $I_{0}=[pq]$ and $I_{K+1}=[pr]$. For all $k\in\{0,\ldots,K\}$, the geodesics $I_{k},I_{k+1}$ together with a subarc of $[qr]$ bound a triangle $T_{k}$.
@@
    For each triangle $T_{k}$ apply the following. Let $\mathcal{A}_{k}$ denote the collection of polygons of the form $A_{i}\cap T_{k}$ intersecting the interior of $T_{k}$; denote this polygon by $T_{k}^{i}$. By ˜2.2, $T_{k}^{i}$ is a convex polygon. Note that each vertex of each $T_{k}^{i}$ is contained in $I_{k}\cup I_{k+1}$. This implies that every edge of some $T_{k}^{i}$ not contained in $I_{k}$ or $I_{k+1}$ must connect $I_{k}$ to $I_{k+1}$. Enumerate such distinct edges as $\{J_{k}^{j}\}_{j=1}^{J_{k}}$ in increasing order from $[qr]$ to the vertex $p$. Also, denote the subarc of $[qr]$ contained in $T_{k}$ by $J_{k}^{0}$. Note that $J_{k}^{j}$ and $J_{k}^{j+1}$ along with the relevant subarcs of $I_{k}$ and $I_{k+1}$ form either a triangle or quadrilateral. We subdivide each triangle $T_{k}$ as follows. For all $j\in\{0,\ldots,J_{k}-1\}$, if $J_{k}^{j}$ and $J_{k}^{j+1}$ do not share an endpoint, then connect an endpoint of $J_{k}^{j}$ to an endpoint of $J_{k}^{j+1}$ by a geodesic, forming two triangles. In this manner we obtain a triangulation $\mathcal{T}$ of $T_{k}$. The final product is shown in Figure˜2(c).
@@
It follows from construction that $\mathcal{T}$ is a vertex-edge triangulation of $T$. Indeed, $\mathcal{T}$ is obtained first by drawing geodesics from $p$ to the opposite edge to form the subtriangles $T_{k}$. Then each $T_{k}$ is the union of triangles satisfying the requirements of ˜4.3. We conclude that $\mathcal{T}$ is vertex-edge.
Finally, we claim that, for each $A\in\mathcal{A}$, the collection $\mathcal{T}_{A}$ can be obtained by vertex-edge subdivision of $A$. Observe that the geodesics $\{I_{k}\}_{k=1}^{K}$ divide $A$ into the non-overlapping union of triangles and quadrilaterals satisfying the requirements of ˜4.3. We now apply this lemma to conclude that $\mathcal{T}_{A}$ is a vertex-edge triangulation of $A$. ∎
@@ 4.2. Subdividing triangles
We note that statement (1) is standard; see Lemma III.9 in [1] for an example of a more general statement.
@@
Next, to prove (2), the main task is to show that the comparison triangles $\bar{T}_{1}$ and $\bar{T}_{2}$ can be fitted into $\bar{T}$ such that $\bar{T}_{1}$ and $\bar{T}_{2}$ are non-overlapping. Our argument is based on Alexandrov’s Lemma; see [7, I.2.16]. Consider the triangle $\bar{T}_{1}$ arranged to share the side $\bar{p}\bar{q}]\subset\bar{T}$ and have overlapping interior. Similarly, the triangle $\bar{T}_{2}$ is arranged to share the side $\bar{p}\bar{r}]\subset\bar{T}$ and have overlapping interior. Let $\bar{y}$ denote the third vertex of $\bar{T}_{1}$ and let $\bar{z}$ denote the third vertex of $\bar{T}_{2}$. The situation is summarized in Figure˜4.3.
@@
Since $\bar{\beta}_{1}\leq\bar{\beta}$ and $\alpha_{1}\leq\alpha$, we see that $\bar{y}$ is contained in $\bar{T}$. Likewise, $\bar{z}$ is also contained in $\bar{T}$. This implies that the triangles $\bar{T}_{1}$ and $\bar{T}_{2}$ are entirely contained in $\bar{T}$. Moreover, since $\bar{\alpha}\geq\bar{\alpha}_{1}+\bar{\alpha}_{2}$, the segments $\bar{p}\bar{r}]$, $\bar{p}\bar{z}]$, $\bar{p}\bar{y}]$, $\bar{p}\bar{q}]$ emanate from $\bar{p}$ counterclockwise in this order, with possibly $\bar{p}\bar{z}]$, $\bar{p}\bar{y}]$ coinciding. From this it follows that $\bar{T}_{1}$ and $\bar{T}_{2}$ do not overlap. We conclude that $|\bar{T}_{1}|+|\bar{T}_{2}|\leq|\bar{T}|$. ∎
@@ 5. $\operatorname{CAT}(\kappa)$ surfaces have bounded curvature
We now proceed with the proof of ˜1.1. Let $X$ be a $\operatorname{CAT}(\kappa)$ surface and let $x\in X$. The first step is to find a neighborhood of $x$ for which we can verify the condition of bounded curvature. Let $r\in(0,D_{\kappa}/2)$ be sufficiently small so that $B(x,r)$ is a $\operatorname{CAT}(\kappa)$ space. Since $r<D_{\kappa}/2$, the set $B(x,r)$ is necessarily convex and contractible in $X$, and in particular is a topological open disk. We assume without loss of generality that $\kappa=1$. Use ˜2.4 to find a triangulation $\mathcal{S}$ of $B(x,r)$ such that each $T\in\mathcal{S}$ satisfies $d(T,\partial B(x,r))>\text{diam}(T)$. Let $S_{1},\ldots,S_{n}$ be the collection of triangles in $\mathcal{S}$ that contain $x$. Then $S=\bigcup_{i=1}^{n}S_{i}$ is a closed topological disk such that $x\in S^{\circ}$. To see this, observe from the convexity of each triangle and uniqueness of geodesics that any two distinct triangles $S_{i}$, $S_{j}$ intersect in a possibly degenerate arc containing the element $x$. We claim that $S^{\circ}$ is the neighborhood of $x$ as required by the definition of surface of bounded curvature.
Let $\mathcal{T}=\{T_{j}\}_{j=1}^{m}$ be an arbitrary collection of simple non-overlapping triangles in $S^{\circ}$. Our goal is to find a uniform upper bound on $\delta(\mathcal{T})$. We apply ˜4.4 to find for each $T_{j}\in\mathcal{T}$ a triangulation $\mathcal{T}_{j}$ of $T_{j}$ that is a vertex-edge refinement of the collection of polygons $S_{i}\cap T_{j}$, $i=1,\ldots,n$, intersecting the interior of $T_{j}$ with respect to $T_{j}$. Note that ˜2.2 implies that each set $S_{i}\cap T_{j}$ is in fact a polygon provided that its interior is non-empty. Observe from ˜4.5 that $\delta(T_{j})\leq\delta(\mathcal{T}_{j})$.
Next, for each triangle $S_{i}$, let $\mathcal{S}_{i}$ be the set of triangles in $\bigcup_{j}\mathcal{T}_{j}$ contained in $S_{i}$. We apply ˜4.4 a second time to $\mathcal{S}_{i}$ to find a vertex-edge refinement of $\mathcal{S}_{i}$ with respect to $S_{i}$. Denote this by $\mathcal{S}_{i}^{\prime}$. Suppose that $T\in\mathcal{T}_{j}$ and contained in $S_{i}$. Let $\mathcal{S}_{i}^{\prime}(T)$ denote the subset of $\mathcal{S}_{i}^{\prime}$ contained in $T$. Then, since $\mathcal{S}_{i}^{\prime}(T)$ is a vertex-edge triangulation of $T$, we have hence $\delta(T)\leq\delta(\mathcal{S}_{i}^{\prime}(T))$ by ˜4.5. This gives
@@
Here, $\bar{T}_{1}^{\prime}$ is the model triangle in $M_{1}^{2}$ corresponding to $T^{\prime}$. Similarly, for each triangle $S_{i}$ we let $\bar{S}_{i}$ denote the model triangle of $S_{i}$ in $M_{1}^{2}$. Since $\mathcal{S}_{i}^{\prime}$ is a vertex-edge refinement with respect to $S_{i}$, ˜4.5 gives
@@
6. Excess-curvature inequality
In this section, we give a proof of ˜1.2 relating the excess and total curvature of an arbitrary Jordan triangle in a $\operatorname{CAT}(\kappa)$ surface. First we verify ˜1.3, that the Hausdorff area of a surface $X$ of bounded curvature without cusp points coincides with the Alexandrov area (see Section˜3.4). This is an application of a theorem of Burago [9] that the metric tangent at each point is a cone over a circle; see also Theorem 9.10 in [36]. Recall that, for a Borel set $A\subset X$, we use $|A|$ to denote the Hausdorff area (i.e., $2$-measure) and $\text{Area}(A)$ to denote Alexandrov area.
Proof of ˜1.3.
We must show that $|A|=\text{Area}(A)$ for an arbitrary Borel set $A$. Since both Hausdorff area and Alexandrov area are determined by their values on polygons, without loss of generality we may assume that $A$ is a polygon. Let $A\subset X$ be a polygon. We can consider $A$ as its own metric space with the induced length metric; this does not affect either notion of area. Let $\varepsilon>0$ be given. For each $x\in X$, let $B_{x}$ be an open metric ball at $x$ with the property of being $(1+\varepsilon)$-bi-Lipschitz equivalent to a ball in a Euclidean cone over a circle, and let $\mathcal{U}=\{B_{x}:x\in A\}$. Let $\mathcal{T}$ be a triangulation of $A$ subordinate to the cover $\mathcal{U}$ and having mesh $\varepsilon$ given by ˜2.4. Define a polyhedral surface $A_{\varepsilon}$ by replacing each triangle $T\in\mathcal{T}$ with the corresponding model triangle $\bar{T}_{0}$. It is clear that $|A_{\varepsilon}|=\text{Area}(A_{\varepsilon})$. Moreover, we have that each triangle $T\in\mathcal{T}$ is $(1+\varepsilon)$-bi-Lipschitz equivalent to the triangle $\bar{T}_{0}$, which implies that
We now proceed with the proof of ˜1.2.
Proof of ˜1.2.
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface and let $T\subset X$ be a Jordan triangle with angles $\alpha,\beta,\gamma$. We may consider $T$ as its own subspace of $X$, having the induced length metric $d_{T}$. Observe that $d_{T}(x,y)\geq d(x,y)$ for all $x,y\in T$, with equality if $x,y$ belong to the same edge of $T$. We conclude that passing to the metric $d_{T}$ does not decrease the angles $\alpha,\beta,\gamma$ and hence does not decrease the excess $\delta(T)$.
We now construct a sequence of model polyhedral surfaces $T_{n}$ converging uniformly to $T$, where each face is a triangle of constant curvature $\kappa$ of diameter at most $\varepsilon_{n}$. Note that such a sequence can be found by invoking the general theory of surfaces of bounded curvature. See Theorem III.10 of [1] for the $\kappa=0$ case; the generalization to arbitrary $\kappa$ is a simple modification. However, for completeness, we provide details for this step.
Fix a decreasing sequence $(\varepsilon_{n})$ of positive real numbers limiting to 0. Assume further that $\varepsilon_{1}$ is sufficiently small that for each point $x\in T$ the ball $B(x,2\varepsilon_{1})$ is $\operatorname{CAT}(\kappa)$. Apply ˜2.4 (with arbitrary choice of open cover $\mathcal{U}$) to obtain a sequence of triangulations $\mathcal{T}_{n}$ of $T$ corresponding to the parameters $\varepsilon_{n}$.